OpenStudy (anonymous):

Translate this problem to an equation and solve. If a number is subtracted from its square, the result is 2. Find the number.

OpenStudy (rajat97):

this can be done like this let the number be x so as the condition says the number 'x' subtracted from its square ' x^2 ' gives two so, we get the equation (x^2)-x=2 then , we get (x^2)-x-2=0 then by solving the quadratic equation, you get the answer as x=2 or -1 so the answer is that x is 2 or -1 hope this helps you

OpenStudy (anonymous):

yes thanks for the help always gratful

OpenStudy (rajat97):

you're welcome and thanks for the medal:)

OpenStudy (anonymous):

Translate this problem to an equation and solve. The length of a rectangle is 8 inches greater than its width. Find the width of the rectangle if its area is 105 square inches.

OpenStudy (anonymous):

your welcome

OpenStudy (rajat97):

just a sec

OpenStudy (rajat97):

i'm goin out for a minute please wait

OpenStudy (rajat97):

i'm back

OpenStudy (anonymous):

k

OpenStudy (rajat97):

let the width be w so according to this problem, the length is 8 inches more than the width so its length=w+8 and its area is given 105 inc.sqr. we know that area is (length)(width) which is given 105 so (w+8)(w)=105 so we get the equation as (w^2)+8w-105=0 so by solving this equation, we get w=15 or -10 any side cannot have its measure as a negative value so we take the width=w=15 and thus, the length=w+8=23inches :)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Translate this problem to an equation and solve. Twelve more than the square of a number is seven times the number. Find the number.

OpenStudy (rajat97):

thanks for becoming my fan

OpenStudy (rajat97):

okay now as we proceed in every question, we'll do the same here so let the number be x this question says that 12 more than the square of this number 'x^2' is 7 times the number 'x' so we get the equation 12(x^2)=7x dividing both the sides by x, we get 12x=7 so x=7/12 remember that if you cancel an x on botht the sides by division, you should even get the value of x as 0 if you dont believe, you can even put 0 in the place of x in the equation here and you'll get 0=0

OpenStudy (rajat97):

thanks for the medal@bibbi

OpenStudy (anonymous):

your welcome